# mimic-video: Video-Action Models for Generalizable Robot Control Beyond VLAs

### v1

@@ IV-F Training
The second stage focuses on learning the action decoder $\pi_{\theta}$ while keeping the video backbone frozen. We train the decoder from scratch to regress the action flow field, conditioned on video representations $\mathbf{h}_{\tau_{v}}$ extracted from the frozen backbone. Crucially, to ensure robustness to varying noise levels during inference, we sample independent flow times $\tau_{v}$ (for video) and $\tau_{a}$ (for action) during each training iteration, as detailed in Algorithm 2. We employ a logit-normal distribution for $\mathcal{T}_{v}$ matching the video pretraining, and $\mathcal{T}_{a}(\tau_{a})\propto\sqrt{\tau_{a}-0.001}$ for actions following [3]. This decoupled training scheme renders our approach significantly more sample-efficient and faster to converge than comparable VLA baselines (see Sec. V-B).
@@ V Experiments
Figure 2: We train and evaluate mimic-video on a real-world bimanual robot setup with Franka Emika Panda robot arms and mimic 16-DoF dexterous humanoid hands. We execute a real-world evaluation on the bimanual setup with two tasks: package sorting and pick and place of a measuring tape into a box. For each action chunk, mimic-video generates a partially denoised video plan, and then executes the actions on the real robot.
@@ V-C Trade-offs between Video Fidelity and Action Performance
Figure 5: Policy success rate vs video noise level ($\tau_{v}$). We sweep the video flow time $\tau_{v}$ across the SIMPLER-Bridge environments. Performance peaks at an intermediate noise level, confirming that high-fidelity video reconstruction is not required for performant robot policies.
Counterintuitively, we find that the best autonomous policy performance in our SIMPLER experiments is achieved at the highest noise level $\tau_{v}=1$ (Fig. 5). Theoretically, as $\tau_{v}$ progresses from $1$ (pure noise) to 0 (full reconstruction), the underlying video signal grows, and the mutual information $I(\mathbf{z}^{\tau_{v}}_{\text{future}};\mathbf{A}^{0})$ between the future video latent and future actions increases. However, consistent with the case study in Sec. III, we hypothesize that imperfect video generation introduces artifacts. Consequently, fully denoised video latents may diverge from the training distribution, presenting out-of-distribution conditioning to the action decoder. To isolate the effect of these generation errors, we perform an additional sweep of $\tau_{v}$ where we condition the action decoder on “noisy” ground-truth video latents, $\mathbf{z}_{\tau_{v}}^{\text{future}}$, computed via Eq. 1. We report the resulting action reconstruction MSE on a held-out validation set of BridgeDataV2 in Fig. 6. We observe that the lowest action reconstruction error is achieved at an intermediate flow time of $\tau_{v}\approx 0.4$, corresponding to with the perfect rollout performance observed in our Case Study (Sec. III).
@@
Interestingly, action prediction error increases sharply as we move from this optimum towards $\tau_{v}=0$ (full reconstruction). We attribute this to the nature of the conditioning signal: while the video latents themselves contain more information at lower noise, the intermediate video model representations —from which the action decoder reconstructs actions— may exhibit distinct, non-trivial behavior. We provide a detailed discussion of these mechanisms in Appendix -D. This observation yields a significant practical advantage: operating at $\tau_{v}=1$ requires only a single forward pass of the video backbone to generate conditioning features, resulting in both the highest average performance and the fastest inference speed.
@@ Acknowledgments
This work was supported under project ID #36 as part of the Swiss AI Initiative, through a grant from the ETH Domain and computational resources provided by the Swiss National Supercomputing Centre (CSCS) under the Alps infrastructure. mimic robotics provided experimental infrastructure, real-world robot platforms and additional compute resources. Primary work by the lead authors was performed during their internships at mimic robotics, with continued development supported during their internships at Microsoft.
We thank Stefanos Charalambous, Erik Bauer, German Rodriguez, Sigmund Hennum Høeg, Irvin Totic and Benedict Wüest for their help with the project.

### v2

@@ IV-F Training
The second stage focuses on learning the action decoder $\pi_{\theta}$ while keeping the video backbone frozen. We train the decoder from scratch to regress the action flow field, conditioned on video representations $\mathbf{h}^{\tau_{v}}$ extracted from the frozen backbone. Crucially, to ensure robustness to varying noise levels during inference, we sample independent flow times $\tau_{v}$ (for video) and $\tau_{a}$ (for action) during each training iteration, as detailed in Algorithm 2. We employ a logit-normal distribution for $\mathcal{T}_{v}$ matching the video pretraining and $\mathcal{T}_{a}(\tau_{a})\propto\sqrt{\tau_{a}-0.001}$ for actions following [3]. This decoupled training scheme renders our approach significantly more sample-efficient and faster to converge than comparable VLA baselines (see Sec. V-B).
@@ V Experiments
Figure 2: We train and evaluate mimic-video on a real-world bimanual robot setup with Franka Emika Panda robot arms and mimic 16-DoF dexterous humanoid hands. We execute a real-world evaluation on the bimanual setup with two tasks: package sorting and pick and place of a measuring tape into a box. For each action chunk, mimic-video generates a latent video plan ($\tau_{v}=1$) and then executes the actions on the real robot. We further fully denoise the predicted video for this visualization.
@@ V-C Trade-offs between Video Fidelity and Action Performance
Figure 5: Policy success rate across the SIMPLER-Bridge environments vs video flow time ($\tau_{v}$, logit-scaled). Performance peaks at an intermediate noise level, confirming that high-fidelity video reconstruction is not required for performant robot policies.
@@
Counterintuitively, we find that the best autonomous policy performance in our SIMPLER experiments is achieved at the highest flow time $\tau_{v}=1$. Theoretically, as $\tau_{v}$ progresses from $1$ (pure noise) to 0 (full reconstruction), the underlying video signal grows, and the mutual information $I(\mathbf{z}^{\tau_{v}}_{\text{future}};\mathbf{A}^{0})$ between the future video latent and future actions increases. However, consistent with the case study in Sec. III, we hypothesize that imperfect video generation introduces artifacts. Consequently, fully denoised video latents may diverge from the training distribution, presenting out-of-distribution conditioning to the action decoder. To isolate the effect of these generation errors, we perform an additional sweep of $\tau_{v}$ where we condition the action decoder on “noisy” ground-truth video latents, $\mathbf{z}_{\tau_{v}}^{\text{future}}$, computed via Eq. 1. We report the resulting action reconstruction MSE on a held-out validation set of BridgeDataV2 in Fig. 6. We observe that the lowest action reconstruction error is achieved at an intermediate flow time of $\tau_{v}\approx 0.4$, corresponding to the perfect rollout performance observed in our Case Study (Sec. III).
Interestingly, action prediction error increases sharply as we move from this optimum towards $\tau_{v}=0$ (full reconstruction). We attribute this to the nature of the conditioning signal: while the video latents themselves contain more information at lower noise, the intermediate video model representations—from which the action decoder reconstructs actions—may exhibit distinct, non-trivial behavior. We provide a detailed discussion of these mechanisms in Appendix -E. This observation yields a significant practical advantage: operating at $\tau_{v}=1$ requires only a single forward pass of the video backbone to generate conditioning features, resulting in both the highest average performance and the fastest inference speed.
@@ Acknowledgments
This work was supported under project ID #36 as part of the Swiss AI Initiative, through a grant from the ETH Domain and computational resources provided by the Swiss National Supercomputing Centre (CSCS) under the Alps infrastructure. We thank mimic robotics for providing experimental infrastructure, real-world robot platforms and additional compute resources. Primary work by the lead authors was performed during their internships at mimic robotics, with continued development supported during their internships at Microsoft.
We thank Benjamin Estermann, Stefanos Charalambous, Erik Bauer, German Rodriguez, Sigmund Hennum Høeg, Irvin Totic and Benedict Wüest for their help with the project.